# Privy: Envisioning and Mitigating Privacy Risks for Consumer-facing AI Product Concepts

### v1

@@ 7.1. Empowering AI practitioners to address privacy risks
Prior work has highlighted how the lack of practical tooling is a key barrier preventing practitioners from systematically addressing privacy risks in AI products (Lee et al., 2024a). To address this gap, we developed Privy, a tool grounded in the needs and insights of AI and privacy practitioners (Section 3), to support the identification and mitigation of privacy risks in novel AI product concepts. Our findings show that both Privy-LLM and Privy-Template enabled non-experts to produce privacy impact assessments rated highly by privacy experts (Section 6.1) and perceived as useful and usable by practitioners themselves (Section 6.2.2). More broadly, Privy empowers practitioners to envision and mitigate privacy risks — addressing challenges of awareness, motivation, and ability that often impede privacy work (Section 6.3.2).
From Privy’s development and evaluation, we distill three key shifts that future tools should aim to support — shifts that help non-privacy experts meaningfully engage with privacy work.
@@
Our findings reaffirm that practitioners’ privacy knowledge underpins their awareness and ability to address risks. Yet, echoing prior work (Lee et al., 2024a), participants described company-provided trainings as overly generic, rarely tailored to AI, and often disconnected from their specific product contexts. Privy, by contrast, scaffolded more rigorous and product-relevant engagement with privacy (Section 6.3.2). Practitioners reported that using Privy heightened their awareness of privacy issues and inspired them to seek additional resources when developing mitigation plans (Section 6.3.3). Many grounded their responses in industry best practices and expressed a desire for even more examples aligned with their products (Sections 6.1.2 and 6.3.3).
@@
Practitioners often face organizational resistance when privacy work appears to be misaligned with product team goals — an issue surfaced in both our findings (Section 6.3.3) and prior work (Lee et al., 2024a). Our findings suggest that a promising path forward is to make privacy work inherently collaborative, by involving non-privacy expert stakeholders in privacy discussions and decision-making early in product development (Sections 6.3.2 and 6.3.3). Beyond collaboration, our findings indicate that tools like Privy can help practitioners move past compliance-driven tasks toward becoming privacy champions within their teams (Tahaei et al., 2021; Lee et al., 2024a). With Privy, participants described feeling ownership and motivation to invest more effort in privacy work, rather than treating it as just another worksheet (Section 6.3.2). Importantly, many produced privacy impact assessments of sufficient quality to meaningfully engage privacy experts (Sections 6.1.1 and 6.1.2) — an early signal that tools like Privy can re-balance privacy decision-making power: equipping every practitioner to confidently initiate and sustain privacy conversations around AI product concepts.
Future tools should continue to support these motivational shifts — helping practitioners advocate for privacy internally. Beyond generating assessments, systems could scaffold how findings are shared: guiding practitioners in presenting risks in team meetings or summarizing them for leadership. In doing so, Privy-like tools could foster a culture where practitioners not only participate in privacy work, but take a leading role.
@@
Consumer-facing AI products are often built around a blunt tradeoff: users gain full access to services in exchange for relinquishing private data (Manheim and Kaplan, 2019; Helbing et al., 2017). Yet our findings suggest that practitioners envision more nuanced approaches — enabling end-users to tailor AI functionality based on their privacy preferences, such as limiting specific capabilities. Participants also acknowledged the challenges of this approach, including ensuring product reliability across varying data availability (Sections 6.1.2 and  6.3.1). When supported by tools like Privy, practitioners were able to generate solutions that moved beyond company-centric (often incorrect) assumptions, such as “users do not care about their privacy settings” (McClain et al., 2023; Solove, 2007), toward designs that foreground user agency. By scaffolding reflection on privacy from varied end-users’ perspectives, Privy helped practitioners articulate options that empower users to take greater control of their personal data (Section 6.1.2).
Future work can build on this by developing user-defined, privacy-preserving AI products, where individuals actively configure the utility–privacy balance of their experience. Privy can be extended to guide practitioners in designing such configuration mechanisms, ensuring that tradeoffs are not imposed unilaterally but negotiated through interfaces that respect user choice.
@@ 7.2. Generative AI-powered privacy risk envisioning and mitigation
In this study, we examined how generative AI (GenAI) can support practitioners in envisioning and mitigating privacy risks (RQ2). Our results are promising: augmenting Privy with an LLM (Privy-LLM) enhanced its effectiveness in helping practitioners identify and mitigate privacy risks (Section 6.2.1). Our findings also shed light on a broader design space for integrating GenAI — not only to generate higher-quality privacy impact assessments, but also to enable practitioners to engage more effectively and critically in the assessment process itself (Section 6.3.3).
Building on these insights, we articulate three design objectives for embedding GenAI-powered tools into non-expert practitioners’ workflows to better support their privacy risk envisioning and mitigation practices.
@@
Privacy education is increasingly critical in the AI era, where emerging capabilities often introduce novel and unforeseen privacy harms (Lee et al., 2024b). Our findings suggest that GenAI-powered tools like Privy-LLM can act as privacy coaches, training practitioners as they work — helping them reason through risks and understand key concepts in context (Sections 6.2.3 and 6.3.2).
Rather than relying on static training, these tools can introduce foundational frameworks (e.g., the AI privacy taxonomy (Lee et al., 2024b)) through product-specific examples, reinforcing concepts within the practitioner’s workflow. Extending our earlier notion of product-driven privacy case studies (Section 7.1), GenAI can dynamically surface relevant examples, case studies, and resources as practitioners progress through assessments. An added benefit is adaptability: unlike fixed training modules, GenAI-powered tools can keep privacy guidance current — reflecting rapidly shifting technological and regulatory landscapes. This positions GenAI not just as a knowledge source, but as an ever-learning privacy mentor embedded in practice.
@@
One of GenAI’s most immediate contributions to privacy work is its role as a co-pilot — a thought partner that enhances both the quality and efficiency of practitioners’ privacy impact assessment. Our findings show that GenAI helped participants think beyond their blind spots and speed up tasks like pre-populating risk descriptions — features they found especially valuable (Section 6.2).
As co-pilots, GenAI-powered tools can enable richer forms of human–AI collaboration in privacy risk envisioning and mitigation. Future directions might include comparing risks across alternative implementations or simulating end-user behaviors to refine product concepts. These affordances open up opportunities for design iteration and exploration that would be difficult to achieve manually. Looking ahead, research should go beyond evaluating assessment quality to examine how GenAI-augmented workflows shape actual product decision-making.
@@
While participants valued GenAI-generated content, they also emphasized the need to maintain agency and stay critically engaged in their privacy work (Sections 6.2.3 and 6.3.3). This aligns with prior work showing that critical engagement is heightened when practitioners have confidence in their own work (Lee et al., 2025) — a finding echoed in our study as our participants expressed high self-efficacy when using Privy-LLM (Section 6.1).
@@ 7.3. Limitations
Our work has several limitations. In the evaluation study, participants assessed two researcher-provided AI product concepts rather than their own. This decision enabled a controlled, between-subject comparison of Privy-LLM and Privy-Template (RQ2) and ensured feasibility within the 90-minute session. Pilot tests showed participants could complete only one privacy impact assessment in that time frame. While we ensured participants understood the assigned concepts, their engagement with privacy risks may differ when working on their own products.
Time constraints also shaped our study design. We allocated 25 minutes for risk identification and 15 minutes for mitigation — enough to complete the activities, but shorter than real-world, iterative development cycles. As such, our findings reflect a snapshot of tool use under structured conditions. Additionally, since each participant used Privy only once, we cannot assess long-term adoption or evolving practices.
Finally, our participant pool was limited to practitioners based in the US and the UK. Although our 35 participants spanned diverse roles and organizations, our findings may not fully generalize across global contexts.
@@ 8. Conclusion
Privy is a novel privacy risk-envisioning tool that helps AI practitioners identify and mitigate the most relevant and severe privacy risks posed by their AI product concepts. In a study with 24 practitioners, complemented by evaluations from 13 privacy experts, we showed that Privy is both useful and usable, and enables practitioners to generate high-quality privacy impact assessments. Through grounding risk envisioning and mitigation in product-specific AI capabilities & requirements, likely use-cases, and likely to be impacted stakeholders, Privy ensures privacy impact assessments are both significant and contextual. Our findings further show that generative AI can enhance the risk envisioning and mitigating process: participants used LLM-generated content to surface overlooked risks, validate their reasoning, and expand mitigation strategies. This work opens new directions for designing tools that empower AI practitioners to meaningfully engage with privacy, and for critically shaping the roles generative AI systems can play in supporting privacy-preserving AI innovation.
Appendix A Privy-LLM Privacy Risk Summarizer Interface
@@
Appendix B Privy-Template Interface
Figure 10. The complete worksheet of Privy-Template.
Figure 11. Reference of Privy-Template: AI privacy taxonomy (Lee et al., 2024b).
@@ D.3. Codebook for the Evaluation Study Qualitative Analysis
D.3.1. RQ1: How and to what extent do Privy-LLM and Privy-Template help practitioners identify and mitigate privacy risks with AI product concepts?

### v2

@@ 7.1. Empowering AI practitioners to address privacy risks
Prior work has highlighted how the lack of practical tools is a key barrier preventing practitioners from systematically addressing privacy risks in AI products (Lee et al., 2024a). To address this gap, we developed Privy, grounded in the needs and insights of AI and privacy practitioners (Section 3), to support the identification and mitigation of privacy risks in AI product concepts. Our findings show that both PrivyLLM and PrivyTemplate enabled non-privacy-expert practitioners to produce privacy impact assessments rated highly by privacy experts (Section 6.1) and perceived as useful and usable by practitioners themselves (Section 6.2.2). More broadly, Privy empowers practitioners to envision and mitigate privacy risks — addressing challenges of awareness, motivation, and ability that often impede privacy work (Section 6.3.2).
From Privy’s development and evaluation, we distilled three key shifts that future tools should aim to help non-privacy-expert practitioners meaningfully engage with privacy work in the age of AI.
@@
Our findings reaffirm that practitioners’ privacy knowledge underpins their awareness and ability to address risks. Yet, echoing prior work (Lee et al., 2024a), participants described industrial training as overly generic, rarely tailored to AI, and disconnected from their specific product contexts. Privy, by contrast, scaffolded more rigorous and product-relevant engagement with privacy (Section 6.3.2). Practitioners reported that using Privy heightened their awareness of privacy issues and inspired them to seek additional resources when developing mitigation plans (Section 6.3.3). Many grounded their responses in industry best practices and expressed a desire for more relevant examples (Sections 6.1.2 and 6.3.3).
@@
Practitioners often face organizational resistance when privacy work appears to be misaligned with product team goals — an issue surfaced in both our findings (Section 6.3.3) and prior work (Lee et al., 2024a). Our findings suggest that a promising path forward is to make privacy work inherently collaborative, by involving non-privacy-expert stakeholders in privacy discussions and decision-making early in product development (Sections 6.3.2 and 6.3.3). Beyond collaboration, our findings indicate that tools like Privy can help practitioners move past compliance-driven tasks toward becoming privacy champions within their teams (Tahaei et al., 2021; Lee et al., 2024a). With Privy, participants described feeling ownership and motivation to invest more effort in privacy work, rather than treating it as just another worksheet (Section 6.3.2). Importantly, many produced privacy impact assessments of sufficient quality to engage privacy experts in conversations around product-specific privacy risks (Sections 6.1.1 and 6.1.2) — an early signal that tools like Privy can re-balance privacy decision-making power.
Future tools should continue to support these motivational shifts — helping practitioners advocate for privacy internally. Beyond generating assessments, systems could scaffold how findings are shared: presentations in team meetings or summaries for leadership. In doing so, Privy-like tools could foster a culture where practitioners not only participate in privacy work, but take a leading role.
@@
Consumer-facing AI products are often built around a tradeoff: users gain full access to services in exchange for relinquishing private data (K. M. Manheim and L. Kaplan (2019); 20). Yet our findings suggest that practitioners envision more nuanced approaches — enabling end-users to tailor AI functionality based on their privacy preferences, such as limiting specific capabilities. Participants also acknowledged the challenges of this approach, including ensuring product reliability across varying data availability (Sections 6.1.2 and  6.3.1). When supported by tools like Privy, practitioners were able to generate solutions that moved beyond company-centric (often incorrect) assumptions, such as “users do not care about their privacy settings” (McClain et al., 2023; Solove, 2007), toward designs that foreground user agency. By scaffolding reflection on privacy from varied end-users’ perspectives, Privy helped practitioners articulate options that empower users to take greater control of their personal data (Section 6.1.2).
Future work can build on this success by developing user-defined, privacy-preserving AI products, where individuals actively configure the utility–privacy balance of their experience. Privy can be extended to guide practitioners in designing such configuration mechanisms, ensuring that tradeoffs are not imposed unilaterally but negotiated through interfaces that respect user choice.
@@ 7.2. Generative AI-powered privacy risk envisioning and mitigation
In this study, we examined how generative AI (GenAI) can support practitioners in envisioning and mitigating privacy risks (RQ2). Our results are promising: augmenting Privy with an LLM (PrivyLLM) enhanced its effectiveness in helping practitioners identify and mitigate privacy risks (Section 6.2.1). Our findings also showed how integrating GenAI could generate higher-quality privacy impact assessments and enable practitioners to engage critically in the assessment process itself (Section 6.3.3).
Building on these insights, we articulate three design objectives for embedding GenAI-powered tools into non-privacy-expert practitioners’ workflows to better support their privacy risk envisioning and mitigation practices.
@@
Privacy education is increasingly critical in the AI era, where emerging capabilities often introduce novel and unforeseen privacy harms (Lee et al., 2024b). Our findings suggest that GenAI-powered tools like PrivyLLM can act as privacy coaches, training practitioners as they work — helping them reason through risks and understand key concepts in context (Sections 6.2.3 and 6.3.2).
These tools can introduce foundational frameworks (e.g., the AI privacy taxonomy (Lee et al., 2024b)) through product-specific examples, reinforcing concepts within the practitioner’s workflow. Extending our earlier notion of product-driven privacy case studies (Section 7.1), GenAI can dynamically surface relevant examples, case studies, and resources as practitioners progress through assessments. An added benefit is adaptability: unlike fixed training modules, GenAI-powered tools can keep privacy guidance current — reflecting rapidly shifting technological and regulatory landscapes. This positions GenAI not just as a knowledge source, but as an ever-learning privacy coach embedded in practice.
@@
One of GenAI’s most immediate contributions to privacy work is its role as a co-pilot — a thought partner that enhances both the quality and efficiency of privacy impact assessment. Our findings show that GenAI helped participants think beyond their blind spots and sped up tasks like pre-populating risk descriptions — features they found especially valuable (Section 6.2).
As co-pilots, GenAI-powered tools can enable richer forms of human–AI collaboration in privacy risk envisioning and mitigation. Beyond facilitating early-stage discussions and decision-making, these tools could be extended to be more prescriptive in suggesting risks and potential solutions (e.g., engineering best practices), support hypothesis testing, or help teams interrogate design trade-offs. Future systems may compare risks across alternative technical implementations or simulate end-user behaviors — especially privacy impacts on marginalized groups — to prompt practitioners to refine product concepts. Such capabilities may enable forms of design iteration and exploration that are difficult to achieve manually.
Looking ahead, comprehensive evaluation of the quality of AI-assisted assessments remains essential as suggestions become increasingly product-specific and actionable. For example, it would be prudent to create a ground truth benchmark to measure alignment between LLM-generated and expert privacy impact assessments. At the same time, future research should go beyond evaluating assessment quality to examine how GenAI-augmented workflows shape actual product decision-making.
@@
While participants valued GenAI-generated content, they also emphasized the need to maintain agency and stay critically engaged in their privacy work (Sections 6.2.3 and 6.3.3). This observation aligns with prior work showing that critical engagement is heightened when practitioners have confidence in their own work when using GenAI tools (Lee et al., 2025) — a finding echoed in our study as our participants expressed high self-efficacy when using PrivyLLM (Section 6.1).
@@ 7.3. Limitations
Our work has several limitations. In the evaluation study, participants assessed two researcher-provided AI product concepts rather than their own. This decision enabled a controlled, between-subject comparison of PrivyLLM and PrivyTemplate (RQ2) and ensured feasible completion within the 90-minute session. While we ensured participants understood the assigned concepts, their engagement with privacy risks may differ when working on their own products.
Time constraints also shaped our study design. We allocated 25 minutes for risk identification and 15 minutes for mitigation — enough to complete the activities, but shorter than real-world, iterative development cycles. As such, our findings reflect a snapshot of tool use under structured conditions. Future work could involve longitudinal deployments of tools like Privy with real product teams to observe sustained patterns of use and adoption. Such studies may shed light on how Privy shapes privacy-related product decision-making and business practices over time and how it integrates with — or reshapes — existing privacy impact assessment workflows.
Additionally, in real-world deployments, product teams using Privy-like systems could leverage more advanced LLMs and fine-tune underlying models to better align with their domain and organizational context — beyond the few-shot configuration used in our study. As such, the effectiveness of Privy can be further enhanced in practice.
Finally, our participant pool was limited to practitioners based in the US and the UK. Although our 35 participants from industry (across both the formative study and summative evaluation) spanned diverse roles and organizations, our findings may not fully generalize across all work contexts.
@@ 8. Conclusion
Privy is a novel privacy risk-envisioning tool that helps AI practitioners without privacy expertise identify and mitigate the most relevant and severe privacy risks posed by their AI product concepts. In a controlled experiment with 24 non-privacy-expert practitioners, whose Privy-scaffolded outputs were evaluated by 13 independent privacy experts, we showed that Privy is both useful and usable, and enables practitioners to generate high-quality privacy impact assessments. Privy ensures privacy impact assessments are both significant and contextual by grounding risk envisioning and mitigation in product-specific AI capabilities & requirements, plausible use-cases, and stakeholders who are likely to be impacted. Our findings further show that, with thoughtful friction, generative AI can enhance the risk envisioning and mitigation process: participants used LLM-generated content to surface overlooked risks, validate their reasoning, and expand mitigation strategies. This work opens new directions for designing tools that empower AI practitioners to meaningfully engage with privacy, and for critically shaping the roles generative AI can play in supporting privacy-preserving AI innovation.
Acknowledgements.
This work was generously supported, in part, by NSF SaTC grants #2316768 and CyLab Presidential Fellowship. We also thank our anonymous reviewers for being generous with their time and helping us improve the paper.
Appendix A PrivyLLM Privacy Risk Summarizer Interface
@@
Screenshot of Privy’s Privacy Risk Summarizer interface. At the top, a “Privacy Risk Highlight” panel summarizes three risks — exposure, surveillance, and identification — along with distributions for risk relevancy (one high, two medium), severity (three medium), and mitigation strategy confidence (two agree, one neutral). Below, Section 1 presents product information for the AI Meeting Assistant, including example use cases, AI capabilities, and AI requirements. Section 2 lists the three identified privacy risks: exposure (high relevancy, medium severity), surveillance (medium relevancy, medium severity), and identification (medium relevancy, medium severity). Section 3 outlines mitigation strategies for the identified risks.
Appendix B PrivyTemplate Interface
Figure 10. The complete worksheet of PrivyTemplate.
Screenshot of Privy-Template. The form is divided into four sections: header, product information, privacy risks, and mitigation strategies. At the very top, the header includes fields for product name, product purpose, and product data. Product information section prompts users to enter potential intended or unintended use cases, along with the entities or individuals that will use and will benefit from the AI product, its AI capabilities (What can the AI do?), and AI requirements (What data or infrastructure is needed for this AI to work?). Privacy risks section guides users to identify three privacy risks by specifying the risk type and risk description, listing impacted stakeholders, and then rating each risk’s relevancy and severity. Definitions are provided for high, medium, and low levels of relevancy and severity. Finally, the mitigation strategies section asks users to propose strategies for addressing the identified risks. For each risk, users are prompted to think about strategies and how the product can be designed to enhance users’ awareness of the risk, how the product can be designed to enhance users’ motivation to address the risk, and how the product can be designed to provide users with the ability to manage the risk. Users are then asked to rate their confidence about the mitigation strategies: I’m confident in the proposed mitigation strategies for addressing this risk. (Please rate the degree to which you agree with this statement.).
Figure 11. Reference of PrivyTemplate: AI privacy taxonomy (Lee et al., 2024b).
Screenshot of the AI Privacy Risk Taxonomy provided as a reference in Privy-Template. The taxonomy is organized into four main categories: (1) Data Collection Risks: includes Surveillance (watching, listening to, or recording of an individual’s activities); (2) Data Processing Risks: include six types: Identification (linking specific data points to an individual’s identity), Aggregation (combining various pieces of data about a person to make inferences beyond what is explicitly captured in those data), Phrenology & Physiognomy (inferring personality, social, and emotional attributes about an individual from their physical attributes), Secondary Use (the use of personal data collected for one purpose for a different purpose without end-user consent), Exclusion (the failure to provide end-users with notice and control over how their data is being used), and Insecurity (carelessness in protecting collected personal data from leaks and improper access due to faulty data storage and data practices); (3) Data Dissemination Risks: cover four types: Exposure (revealing sensitive private information that people view as deeply primordial that we have been socialized into concealing), Distortion (disseminating false or misleading information about people), (3) disclosure (revealing and improperly sharing data of individuals), Disclosure (revealing and improperly sharing data of individuals), and Increased Accessibility (making it easier for a wider audience of people to access potentially sensitive information); (4) Invasion Risks: include Intrusion (actions that disrupt or disturb an individual’s solitude in physical space).
@@ C.1. Privybeta’s Interfaces
Screenshot of Privy Beta’s risk explorer interface, shown as a node-based workflow. On the left, a gray box displays the product name and description. Below, a green box outlines AI capabilities and requirements. Arrows connect to blue boxes representing use cases, with labels for intended (green) and unintended (red) uses, each tagged with potential beneficiaries or exploiters of the use case. These use cases link to purple boxes identifying impacted stakeholders, such as “individuals whose contact information is accessed without their consent”. Finally, arrows extend to orange boxes that present privacy risks, such as surveillance, insecurity, aggregation, and exposure, each accompanied by a description of how the risk could manifest and affect the impacted stakeholders.
@@
Screenshot of Privy Beta’s risk summarizer interface. At the top, a green header indicates the use case example “Targeted Fundraising Campaigns.” The left panel lists impacted stakeholders and associated risks, such as surveillance, insecurity, disclosure, and aggregation. The right panel provides real-world incident references, each with metadata such as sector, purpose, technology, and links to related articles. Finally, a button on the top right allows users to generate a unique link to share the report.
@@ D.3. Codebook for the Evaluation Study Qualitative Analysis
D.3.1. RQ1: How and to what extent do PrivyLLM and PrivyTemplate help practitioners identify and mitigate privacy risks with AI product concepts?
@@
D.4. Privacy Risk Types Identified by Practitioners and Privacy Experts
Table 8. We found that both PrivyLLM and PrivyTemplate enabled practitioners to identify a broad spectrum of privacy risks across both AI product concepts. Overall, the risk types they surfaced aligned with those identified by the privacy experts. While practitioners identified a slightly more diverse set of risk types using Privy, this variation should be interpreted tentatively given the study’s limited sample size.
AI Meeting Assistant Dynamic Audience Selection
Counts of Each Risk Type
Selected in PIA Reports
PrivyTemplate
PrivyLLM
Privy
Privacy
Experts*
PrivyTemplate
PrivyLLM
Privy
Privacy
Experts*
Surveillance 6 3 9 7 2 6 8 2
Identification 3 1 4 0 3 2 5 3
Aggregation 2 3 5 1 2 0 2 4
Phrenology Physiognomy 0 0 0 0 0 0 0 0
Secondary Use 2 1 3 3 1 0 1 2
Exclusion 0 2 2 4 2 0 2 0
Insecurity 2 1 3 1 1 1 2 1
Exposure 1 3 4 0 1 2 3 1
Distortion 1 2 3 3 1 2 3 2
Disclosure 1 0 1 1 2 2 4 3
Increased Accessibility 0 1 1 1 2 0 2 0
Intrusion 0 0 0 0 0 3 3 0
*Note. During privacy expert evaluation (Section 5.2.3), each privacy expert was asked to select the top three risk types they saw as the most crucial to address for their assigned AI product concept.